OpenStudy (anonymous):

how to find the equation of the vertical asymptote of h(x)=0.02/(6.5x-2.1), and....is the domain all real numbers except -2.1?

OpenStudy (reemii):

the numbers excluded from the domain are those x who make the denominator equal to zero. 6.5x-2.1=0 , solve for x, you'll obtain: x = 2.1/6.5 the asymptote has equation x = 2.1/6.5 the line is vertical (x=...) and it happens at the excluded value (2.1/6.5)

OpenStudy (anonymous):

so domain is all real numbers except 0.02?

OpenStudy (anonymous):

and do i divide 2.1/6.5 for the equation, or is it 2.1/6.5 the solution?

OpenStudy (anonymous):

thank you, by the way-math is not easy for me..

OpenStudy (reemii):

2.1/6.5 = 0.323.. we can call this number, say, \(\alpha\) and the equation is: \(x=\alpha\). and \(\alpha\) is the number excluded from the domain. (the domain is \(\mathbb{R}\setminus \{\alpha\})\).